What is the probability of getting a sum greater than 4 when you roll two dice?
Answer by stanbon(48569)

What is the probability of getting a sum greater than 4 when you roll two dice?
------------
Total # of sums = 36.
# of sums less than 5: 11,12,13,21,22,31 = 6
# of sums greater than 4: 30
---------
P(sum is greater than 4) = 30/36 = 5/6
=======================================
Cheers,
Stan H.